### 10 Apr 2026 v2

Abstract
Software redesign preserves functionality while improving quality attributes, but manual reuse of code and tests is costly and error-prone, especially in cross-repository redesigns. Focusing on static analyzers where cross-repo redesign needs often arise, we conduct a bidirectional study of the ongoing Soot/SootUp redesign case using an action research methodology that combines empirical investigation with validated open-source contributions. Our study reveals: (1) non-linear migration which necessitates bidirectional reuse, (2) deferred reuse via TODOs, (3) neglected test porting, and (4) residual bug propagation during migrations. We identify tracking corresponding code and tests as the key challenge, and address it by retrofitting clone detection to derive code mappings between original and redesigned projects. Guided by semantic reuse patterns derived in our study, we propose Semantic Alignment Heuristics and a scalable hierarchical detection strategy. Evaluations on two redesigned project pairs (Soot/SootUp and FindBugs/SpotBugs) show that our approach achieves an average reduction of 33–99% in likely irrelevant clones at a SAS threshold of 0.5 across all tool results, and improves precision up to 86% on our benchmark of 1,749 samples. Moreover, we contribute to the redesigned projects by submitting five issues and 10 pull requests, of which eight have been merged.
1 Introduction
Software redesign (or rewrite) is a challenging maintenance task that improves non-functional properties (e.g., extensibility) while preserving behavior (Stuurman et al., 2016). During software redesign, developers must balance maintaining core functionality with major structural improvements. Ensuring functional equivalence often requires manually reusing code and tests, which is costly and error-prone.
Why Static Analyzers? Designing static analyzers that are both precise and efficient is inherently challenging, as architectural changes often ripple through program representations, APIs, and performance-critical components (Wei et al., 2018), making redesign unavoidable. Consequently, the needs for software redesign often arise in the cases of static analyzers. For example, Soot (Lam et al., 2011; Vallée-Rai et al., 2000; Park and Jung, 2022; Arzt et al., 2017, 2013, 2014; Zhang et al., 2024b), a widely used static analysis framework was redesigned as SootUp in 2022 to address long-standing architectural limitations (Karakaya et al., 2024), yet the migration remains incomplete after more than three years, requiring concurrent maintenance of both projects. A similar redesign occurred when FindBugs (Hovemeyer, 2015) was succeeded by SpotBugs (HackerNews, 2025). The ongoing Soot/SootUp redesign provides us with a unique opportunity to conduct an action research study of bidirectional code reuse, whereas the completed FindBugs/SpotBugs migration allows us to evaluate the generalizability of our proposed technique.
These redesigns differ fundamentally from commonly studied redesign scenarios—such as mobile application UI migrations that occur within a single repository and are traceable via commit histories (Bessghaier et al., 2022; Pradana and Nuryuliani, 2023; Fuada et al., 2024; Salvatore and Pamuji, 2025). They typically occur within the same repository and are traceable via commit histories (e.g., AntennaPod (AntennaPod, 2025) migrated from Material 2 to Material 3). In contrast, static analyzer redesigns often span multiple repositories, involve substantial architectural changes, and evolve as independent projects. These characteristics complicate reasoning about functional correspondence across versions.
To understand how such redesigns unfold in practice, we conduct a bidirectional study of the ongoing Soot/SootUp redesign, mining redesign-related changes from software artifacts (e.g., issues and pull requests (PRs)) via cross-references between the two projects. This bidirectional perspective is necessary because developers often modify both projects concurrently.
Our study identifies a recurring challenge: developers struggle to track corresponding production and test code as the original and redesigned projects evolve. While reused fragments across the two projects often resemble code clones, only a subset of these clones reflect meaningful functional migration relationships. Distinguishing these relationships proved difficult in practice and hindered reasoning about redesign progress and correctness.
Motivated by this challenge, we explored whether clone detection techniques could be adapted to identify code mappings between the original and redesigned projects, expressed as method pairs. We represent potential mappings as method pairs (Figure 1), and categorize them into genuine clones, code mappings, irrelevant clones, and non-clones, as described below:
Genuine Clone (GC) denotes a method pair ($m_{1}$, $m_{2}$) satisfying the traditional clone definitions (Type-1 to Type-4, see Section 2(Svajlenko and Roy, 2015; Wang et al., 2023);
Code Mapping (CM) refers to a method pair ($m_{1}\in proj_{orig}$, $m_{2}\in proj_{redesign}$) representing a functional migration link where $m_{2}$ is either the direct one-to-one functional counterpart of $m_{1}$ after migration, or transitively mapped to $m_{1}$ via within-project clones.
Irrelevant Clone denotes any genuine clone that is not a code mapping, representing the noise that hinders precise code mapping identification.
Non-clone denotes all method pairs that fail to satisfy the GC criteria.
Figure 1: Relationship between non-clones, genuine clones, code mappings and irrelevant clones in $D$.
We evaluated four representative method-level clone detection techniques (NiCad, DeepSim, CCSTokener, and an LLM-based approach) (Roy and Cordy, 2008; Zhao and Huang, 2018; Wang et al., 2023; Khajezade et al., 2024) on two redesigned project pairs: the ongoing Soot/SootUp redesign and the completed FindBugs/SpotBugs case. Our evaluation revealed three key challenges: (CH1) prior techniques often produce many false positives (FPs), where the identified code pairs share low semantic similarity but being incorrectly marked as clones due to their structural or syntactic similarities (e.g., tokens, AST paths); (CH2) lack of support for distinguishing meaningful code mappings from irrelevant clones. (CH3) limited scalability of LLM-based approaches (Khajezade et al., 2024) due to costly pairwise comparisons.
To address the aforementioned challenges, we enhance prior clone detection techniques with a set of rules that prioritize preserved natural language information (e.g., identifiers, documentation, and comments) across redesigned projects. We call these rules Semantic Alignment Heuristics (SAH) that capture preserved semantic cues (e.g., identifiers, comments) across original and redesigned projects. To address CH1 and CH2, we compute the Semantic Alignment Score (SAS) for each method pair to filter out unlikely or irrelevant pairs, improving precision while maintaining coverage. To support repository-scale analysis for LLM-based approaches (address CH3), we further designed a hierarchical detection pipeline that performs class-level pre-filtering prior to method-level analysis, substantially reducing computational and monetary costs.
Table 1: Empirical Findings on Code Reuse in Redesigned Projects
Findings on Reuse Characteristics (Section 3) Implications
Redesign-related changes are not localized but fragmented across various artifacts (commits, PRs, issues, documentation), posing challenges for manual tracking. Keyword and label searches are insufficient to recover these links at scale, necessitating automated tools that explicitly track cross-project change relationships.
Redesign is a non-linear migration where parallel maintenance requires bidirectional reuse. Redesign focuses on forward porting of production code, while test reuse remains limited due to significant API changes. Backporting tests can catch residual bugs and unblock PRs but rarely occurs in practice. Redesign projects should adopt a bidirectional porting strategy. Automated tools must bridge the porting gap through bidirectional test migration: forward porting reduces redundant efforts for verification integrity, while backporting leverages tests to resolve legacy defects.
Reuse in redesign adapts to new contexts, lowering textual similarity while preserving semantics through specific patterns: (1) customized renaming, (2) similar identifiers and signatures, and (3) similar comments. These patterns serve as semantic anchors that bridge the textual gap. They help retrofit existing clone detectors to accurately establish code mappings despite extensive structural redesign.
Findings on Clone Detector Efficacy (Section 5) Implications
Test harness within test code leads to more irrelevant clones than in production code. Considering identifier name similarity may help filter this noise in test code mappings.
Manual analysis of code mappings revealed three instances of porting-related bugs with inconsistent behavior. Automated code mapping could enable systematic analysis of porting-related bugs otherwise buried among candidate clones.
Prior clone detection tools are more effective at identifying production code than test code, but can produce false positive rates up to 87%. Existing clone detectors need to be adapted for more effective code mapping detection by enhancing false positive filtering.
Findings on Proposed Approach (Section 5) Implications
Our rules improve clone detection: pre-filtering prunes over 98% of method pairs for GPT-based detectors, enabling repository-scale analysis , while post-filtering reduces false positives, boosting precision by up to 86% and reducing manual inspection by 33–99%. Identifier matching emerges as the strongest semantic signal for code mapping. This suggests future work could leverage such stable signals to guide mappings for changed parts, supporting automated code adaptation and helping developers accelerate porting in redesigned projects.
Similar Identifier Name Matching is the most effective rule, while other rules enhance similarity measurement using different method information.
Overall, this paper makes the following contributions:
  • Action Research Study: We conduct the first bidirectional study on software reuse in redesigned projects from the perspective of static analyzers using an action research methodology (Runeson and Höst, 2009; Avison et al., 1999) where we identified issues in the first exploratory stage, followed by the action stage where we brought about changes by making carefully reviewed and validated open-source contributions to address the issues; in the final stage, we reflected upon the code reuse practice. Our study derives eight key findings that characterize code and test reuse in cross-repository redesign (Table 1). Our contributions to the ongoing Soot/SootUp redesign include: reporting four bugs (three fixed), backporting three fixes with tests, and forward porting three validators to the redesigned project.
  • Redesign code-mapping dataset: We contribute a code mapping dataset for redesigned projects, comprising 1,749 method pairs identified by four clone detection techniques across two redesigned project pairs. We manually categorize the pairs as genuine clones (i.e., Type-1 to Type-4 clones), code mappings, or non-clones. This dataset serves as a benchmark for future research on software reuse in real-world redesign contexts, addressing the current lack of such datasets.
  • Redesign-aware clone filtering: Guided by empirical reuse patterns, we propose redesign-aware heuristics that retrofit existing clone detectors to support code reuse during redesign. Our approach leverages redesign-specific renaming, preserved identifiers, and natural-language cues, reducing 33–99% of irrelevant clones and improving precision to 86% across tools. Moreover, by filtering over 98% of method pairs via class-level pre-filtering, our heuristics enable GPT-based detector to scale to repository-level inputs.
2 Background and Tool Selection
2.1 Redesigned Projects
We introduce two redesigned project pairs: Soot/SootUp and FindBugs/SpotBugs, and justify their selection as research targets.
Soot. Soot is a popular static analysis framework widely used to analyze, instrument, and optimize Java and Android applications (Lam et al., 2011; Vallée-Rai et al., 2000; Park and Jung, 2022; Arzt et al., 2017, 2013, 2014). Sponsored by Amazon Web Services, Soot is an open-source project with over 2,000 stars on GitHub.
SootUp. In December 2022, SootUp was released as a new version of Soot with a completely redesigned architecture. The project aims to restructure Soot away from heavy use of singletons, providing a lighter and more easily embeddable library. It also emphasizes parallelization to improve execution speed (Karakaya et al., 2024). Compared to its predecessor Soot, SootUp includes (SootUp, 2024f): (1) improved API without globals/singletons, (2) parallelizable architecture, (3) lazy class loading, (4) fail-early strategy (input validation during object construction), and (5) new source code frontend. Similar to Soot, SootUp is open-source with over 700 stars on GitHub.
FindBugs. FindBugs is an open-source static analyzer for detecting bugs in Java programs (Hovemeyer, 2015) that analyzes Java bytecode.
SpotBugs. SpotBugs is the spiritual successor of FindBugs, carrying on from the point where it left off with the support of its community. In 2016, the lead of FindBugs became inactive, leaving numerous community issues unresolved. This prompted volunteers to create a modernized, more maintainable project for Java. On September 21, 2017, the first official version of SpotBugs v3.1.0 was released. Unlike FindBugs, it supports language features of newer versions of Java.
Selection of redesigned projects. We select Soot/SootUp and FindBugs/SpotBugs as the target project pairs based on their popularity, and open-source status, using versions 4.5.0/1.3.0and 3.0.1/4.5.0 respectively. First, they have been extensively studied in domains such as automated testing, software evolution, and bug detection (Zhang et al., 2023b; Tomassi, 2018; Wang et al., 2022; Lavazza et al., 2020; Zhang et al., 2023a, 2024a). Second, these projects are important and popular because they are developed and widely used by software engineering and programming language researchers to build automated tools (i.e., many projects (Lam et al., 2011; Vallée-Rai et al., 2000; Park and Jung, 2022; Arzt et al., 2017, 2013, 2014) rely on them and need updates to adopt the redesign versions). Third, their open-source status enables a comprehensive analysis of source code, issues, and PRs.
2.2 Code Clone and Detection Tool
Code clones are typically classified into four types (Svajlenko and Roy, 2015; Wang et al., 2023): Type-1 (exact clones) are syntactically identical, ignoring whitespace and comments. Type-2 extend this to include variations in identifier names and literal values. Type-3 (near-miss clones) involves structural differences like added, modified, or removed statements. Though functional similarity is not required, their syntactic resemblance often yields similar semantics. In contrast, Type-4 are syntactically dissimilar but semantically equivalent (e.g., the example in Figure 10).
Selection of Clone Detection Tools. Analyzing code clones is a common approach to studying code reuse. Several code clone detection tools (Jiang et al., 2007; Roy and Cordy, 2008; Sajnani et al., 2016; Zhao and Huang, 2018; Saini et al., 2018; Wang et al., 2018; Gupta and Goyal, 2021; Wu et al., 2022; Wang et al., 2023; Khajezade et al., 2024) have been proposed. Our study focuses on four tools representing diverse algorithms: NiCad (Roy and Cordy, 2008) (text-based), DeepSim (Zhao and Huang, 2018) (graph- and learning-based), CCStokener (Wang et al., 2023) (token-, tree-, and graph-based), and an LLM-based technique (Khajezade et al., 2024). We selected these tools for their strong performance compared to other clone detectors (Wang et al., 2023, 2018; Wu et al., 2022; Saini et al., 2018) and their support for method-level granularity–a commonly used unit for API documentation and unit testing.
3 Reuse Characteristics and Patterns
Figure 2: Overview of Our Bidirectional Study of an Ongoing Redesigned Project
To understand how developers perform code and test reuse, we conducted a bidirectional study of the unique characteristics of reuse between the original and redesigned projects (Soot vs. SootUp). Figure 2 provides an overview of our study. We collect redesign-related data from the GitHub repositories of both the original project Soot and the redesigned project SootUp, since developers maintain both projects in parallel during the migration of core functionalities. We combine keyword and label searches to identify redesign changes scattered around various software artifacts, including source code, pull requests, commits, issues, and documentation. For the direction from SootUp to Soot, we collected Soot-related data from the SootUp repository using keywords “old soot” and “soot equivalent”. Conversely, for the opposite direction, we used keywords “sootup” and “futuresoot”. We selected these keywords as they have been consistently used by developers when referencing the project. We also collected all closed issues (i.e., labeled as bugs and unlabeled) to identify residual bugs inherited from Soot and potential opportunities for synchronizing fixes between the two projects. Two authors performed independent manual analyses and discussed any conflicts to reach a consensus.
Finding 1: In the Soot/SootUp redesign, related changes are not localized to code but are fragmented across various types of software artifacts (e.g., commits, pull requests, issues, and documentation). This fragmentation poses challenges in tracking the changes between the original and the redesigned projects. Although keyword and label searches can partially recover these links, such approaches are insufficient at scale, motivating the need for automated tools that explicitly track cross-project change relationships.
3.1 Analysis from SootUp to Soot
Keyword Search: We collected 22 code instances, 27 PRs, 13 commits, 30 issues and seven documents, covering a five-year span of historical data obtained through keyword search.
Label Search: We examined 256 closed issues, labeled as bugs or unlabeled. Among these, 37 were referenced by commits (16 labeled as bugs, 21 unlabeled). No reusable patch or test was found, mainly because: (1) keyword matching fails to locate the corresponding file or code in Soot to apply the patch in SootUp; (2) the SootUp bug cannot be replicated in Soot.
3.2 Analysis from Soot to SootUp
Keyword Search: The keyword search found two PRs, two commits, three issues, and two documentation entries, but no code instances. Only a closed PR (Soot, 2022) addressing the always-false condition (Section 3.1) related to code reuse. To identify potential test reuse, we matched code and test classes by file name (e.g., “JNopStmt.java” $\rightarrow$ “JNopStmtTest.java”), revealing 35 test classes missing in Soot 4.5.0 but exist in SootUp 1.3.0. These can be potentially backported as the corresponding code classes exist in both versions.
Label Search: We compiled 678 closed issues over 12 years, categorized as bugs or left unlabeled. Among these, 75 issues were referenced by commits (37 bugs, 38 unlabeled). We did not identify portable code for SootUp due to (1) non-reproducible bugs, and (2) fixes already existing in SootUp.
We manually analyzed the constructed dataset and observed several characteristics (C) and patterns (P) of software reuse. For each characteristic, we take corresponding action (Action) with carefully reviewed contributions, and reflected upon the results (Reflection). In the remaining examples in this paper, red lines denote Soot code, whereas green lines represent SootUp code.
[C1] TODO comments are a hidden form of deferred code reuse which may include important validation checks. Among the 22 code instances, 17 contain TODO comments that embed copied code fragments from Soot as placeholders for future adaptation. These TODOs correspond to validation methods, which perform integrity and correctness checks on generated Jimple or other intermediate representation (IR) bodies. Although such validation logic constitutes non-functional requirements, it is critical for ensuring the quality of generated IRs. However, our study revealed SootUp developers tend to defer implementing non-functional requirements (e.g., validation methods), incurring technical debt that may compromise the quality of generated IRs. Action: To reduce this debt, we submitted five PRs to port several validation methods from Soot, including test cases; three have been merged (SootUp, 2024a, 2025a, 2025b). As of the release of SootUp 1.3.0, eight validation methods were adapted where six are Type-3 clones and two are Type-4 clones of Soot counterparts. Reflection: Our analysis reveals that deferred code reuse during redesign often goes beyond copy-paste, requiring semantic reinterpretation to fit the new context. In practice, SootUp developers adapted code in the TODO comments by (1) changing method signatures (parameter and return types); (2) adapting names of variables and invoked methods (e.g., “unit” to “stmt”); (3) adjusting internal API usage (e.g., Soot uses body.getMethod(), while SootUp uses view.getMethod(body.getMethodSignature())); and (4) introducing functional divergence, which may result in Type-4 clones (e.g., Soot’s MethodDeclarationValidator checks valid method signatures, while SootUp additionally enforces that signatures avoid impossible modifier combinations). These changes reduce textual similarity to the original code, making detection more challenging.
1public List<ValidationException> validate(Body body,View view){
2 // TODO: check copied code from old soot
3 /* SootMethod methodRef = body.getMethod(); ... } */
4 return null; }
Figure 3: SootUp method body copied from Soot code as TODO comments for future reuse
In issues, commits, and PRs, we focused on keyword-related descriptions and patches. We also obtained all commits related to code reuse that belong to the collected PRs. Comparing the code in PR patches with the corresponding code in the latest Soot release, we manually identified code clones in 8 PRs. We excluded 19 PRs as they were either irrelevant to code in Soot or contain enormous changes (i.e., difficult to isolate and pinpoint related changes). Within the 8 PRs, we identified characteristic [C2].
[C2] Forward porting commonly occurred in the source code of redesigned projects instead of test code, whereas backporting of code/tests rarely occurred. Inspection of PR patches shows that most porting from Soot to SootUp occurred in source code instead of test code, aligning with previous findings that forked projects rarely benefit from tests created in other forks (Mukelabai et al., 2023). The substantial amount of untested functionalities in Soot highlights missed opportunities for backporting tests. Several factors may limit test reuse: (1) Soot had many commits without tests, indicating inadequate test coverage; (2) significant API changes in SootUp compared to Soot (Karakaya et al., 2024) reduced code similarity, limiting test reuse; (3) SootUp meticulously designed its testing framework and test cases. For example, when porting the DeadAssignmentEliminator class, the internalTransform method in Soot was renamed to interceptBody in SootUp. However, the corresponding test class DeadAssignmentEliminatorTest had completely different tests. In Soot, each test constructs its own input, whereas SootUp uses fewer inputs but more reusable inputs (e.g., several tests share a createBody() helper method (SootUp, 2020b)). Action: To encourage backporting, we contributed by backporting bug-fix-related pull requests. For example, we notice that a PR was previously submitted to Soot but was closed due to missing tests, while SootUp reused the fix and added corresponding tests. To address this, we submitted a PR to backport both the fix and its tests, which was accepted (Soot, 2024b). Reflection: Our experience of backporting reveals that backporting fixes with associated tests may help unblock issues/PRs that were previously closed due to missing tests, which can subsequently help improve the quality of the original project (Soot).
1private void convert() { ... 2- Edge edge = worklist.pollLast(); 3- AbstractInsnNode insn = edge.insn; 4+ BranchedInsnInfo edge =worklist.pollLast(); 5+ AbstractInsnNode insn = edge.getInsn();
Figure 4: Identifier name reuse with type change in SootUp (SootUp, 2021b) compared to Soot
1-protected void internalTransform(Body b, String phaseName, Map<String, String> options) ... 2+public void interceptBody(Nonnull Body.BodyBuilder builder) ... 3// Make a first pass through the statements, noting the statements we must absolutely keep. 4 ... // Stmt is of the form a = a which is useless ... 5 // Remove the dead statements 6- units.retainAll(essential); 7+ for (Stmt stmt : stmts) { ...
Figure 5: Similar inline comments between SootUp (SootUp, 2020b) and Soot
1/** Utility methods for string manipulations commonly used in Soot. */
2 public class StringTools {
3- /**Returns fromString, but with non-isalpha() characters printed
4- as $\unnnn’</code>$.
5- Used by SootClass to generate output. */
6+ /**Returns fromString, but with non-isalpha() characters printed
7+ as $\unnnn’</code>$.*/
8 public static String getEscapedStringOf(String fromString) {...} ...}
Figure 6: Similar Javadoc comments in a SootUp patch (SootUp, 2020a) and the corresponding in Soot (4.5.0)
[C3] Propagation of residual bugs during code migration and detection of new bug in reused code block. Similar to the observation in a prior study (Mondal et al., 2019), our analysis revealed that most code clones in SootUp (with respect to Soot) are Type-3 clones, leading to residual bug propagation. Specifically, we discovered two fixes in SootUp for residual bugs that were initially reported in Soot (SootUp, 2022, 2021a). As Soot had left these bugs unfixed, they were inherited by SootUp during code migration. One bug involved an always-false condition in AsmMethodSource, leading to an unreachable branch. The SootUp PR (SootUp, 2022) referenced a closed Soot PR (Soot, 2022), and its fix was a refined version of that Soot PR, with fewer changes and additional tests. The other bug resulted in a crash during method body retrieval. Action: To resolve the identified residual bugs in Soot, we created two PRs to backport the fixes, along with the associated tests, from SootUp. Both PRs were merged by the Soot’s developer (Soot, 2024a, b) with positive comments such as “Thank you for this effort, much appreciated…”. The test for the second residual bug was ignored in SootUp. While trying to forward-port the test, we found a new bug within a reused code block. The bug is caused by the initialization of a field trapHandler in Soot involved two steps. However, in SootUp, a method call buildTraps() was inserted, referencing the field before its initialization, resulting in a null handler exception. We reported this bug (SootUp, 2024e), and the SootUp’s developer responded quickly by fixing the issue. Reflection: Fixing residual bugs shows that code migration can propagate existing defects and introduce new ones, highlighting the need for careful review of modifications to reused code to preserve functionality.
Finding 2: Our study reveals three key insights on the code reuse in the ongoing Soot/SootUp redesign: (1) Redesign is not always a linear migration, as parallel maintenance of two projects necessitates bidirectional reuse through forward porting and backporting. (2) Redesign mainly focuses on forward porting of production code, while test reuse remains limited despite being theoretically easier, likely due to significant changes in SootUp’s internal APIs and testing framework that hinder direct test reuse. (3) Backporting rarely occurred for both production code and tests. Our study reveals that many Soot classes remain untested despite the existence of corresponding tests in SootUp, indicating missed backporting opportunities that could catch residual bugs and unblock previously closed PRs due to missing tests.
[P1] Customized renaming rules during redesign. Using keyword search, we observed that many identifiers in SootUp were systematically renamed during redesign, introducing textual divergence. These renaming patterns break the direct correspondence between text and semantics, highlighting a common source of semantic-textual mismatch during redesign.
  • To ensure immutability, SootUp uses withers instead of setters (e.g., withName in SootUp maps to setName in Soot(Karakaya et al., 2024).
  • Concept replacement (e.g., View in SootUp maps to the singleton Scene in Soot, BodyInterceptors replaces the concept of BodyTransformer, and Stmt corresponds to Unit. Consequently, methods like getStmt correspond to getUnit).
Based on the matched code, we derived code reuse patterns, illustrated below with examples.
[P2] Similar identifier names and method signatures (8/8). We notice that code in SootUp often uses identical or similar identifier names (e.g., class/field/method/variable names) and method signatures from Soot. For example, Figure 5 shows a code reuse instance in AsmMethodSource::convert, wherein SootUp retains the class name and method signature without alteration. The names of the local variables (e.g., edge, ins) remain unchanged despite changes in their data types. This likely occurs because during redesign, developers tend to restructure the code by adding new classes while the underlying natural language channel remains unchanged (i.e., the variables involved remain the same).
[P3] Similar inline and Javadoc comments (5/8). SootUp developers often reuse comments within methods. Figure 5 shows matching inline comments despite signature and logic changes, aiding code reuse identification. Figure 6 further shows identical class Javadoc but shortened method Javadoc in SootUp, reflecting behavioral or usage changes.
Based on these patterns, we derived rules to improve clone detection in redesigned projects (Table 2). P1 applies regex-based identifier renaming ( Table 3). P2 compares method headers (return type, method name and parameters) and local variables (variable type and name) to identify similar methods as a lightweight alternative to full-body analysis. P3 measures documentation and comment similarities to capture reused descriptions.
Table 2: Rules derived from patterns discovered in our study
Pattern Rule Description
P1 R1: Redesign-aware Renaming Renaming identifiers across scopes based on custom rules in redesign documentation.
P2 R2: Similar Identifier Name Matching Measure similarity in method headers and local variables to identify similar methods.
P3 R3: API Documentation-aware Matching Measure method API documentation similarity to identify semantically similar methods.
R4: Inline Comment-aware Matching Measure inline comment similarities to identify method clones from copy-paste instances.
Finding 3: Code reuse in redesign adapts to the new context, often lowering textual similarity to the original code. We identified three textual patterns that preserve semantics and can enhance using clone detection tools for redesigned projects: (1) customized renaming rules during redesign (e.g., using withers instead of setters), (2) similar identifiers and method signatures despite distinct data types, and (3) similar comments in code clones across projects.
4 Methodology of Redesign-aware Clone Filtering
Figure 7: Overview of our redesign-aware clone filtering workflow that retrofit existing clone detectors.
Building upon our case study, we extend prior clone detectors with the reuse rules in Table 2 to improve code mapping for redesigned projects. Figure 7 shows our conditional hierarchical workflow: for a redesigned project pair, we extract method details while optionally applying class-level pre-filtering for detectors with limited scalability (e.g., LLM-based ones). We then run clone detectors to identify cross-project method clones. For each candidate pair, we preprocess their details and compute a semantic alignment score (SAS). These scores are then used in a method-level post-filtering step to produce the final code mappings.
Data Extraction. The data extraction step parses code files to collect method details for similarity computation, including the containing file and class of each method , method header (return type, name, parameters), local variables, inline comments, and docstrings. We exclude interfaces, abstract methods and constructors, as they lack concrete logic. We also exclude common methods resulting from object-oriented design (e.g., toString() in the Object class) to focus on code mappings between redesigned projects.
4.1 Class-level Pre-Filtering for LLM-based Detectors
To scale LLM-based clone detectors to repository-level inputs, we apply class-level filters. Our goal is to drastically reduce the number of method pairs before feeding them to clone detectors. Following pattern P2, our rationale is that methods from class pairs where their names are of low similarity is unlikely to be correct mappings. We first normalize equivalent concepts in fully qualified class names (R1), then compute class name similarity (R2) and discard pairs with low scores. We only pair methods within each retained class pair, and further discard pairs with a line-count ratio $\geq 2$ or embedding similarity <0.5<0.5.
4.2 Preprocessing
[…]
Table 3: Implemented redesign-aware replacement rules
Scope Target Pattern Matching Replacement Examples
All Soot (Unit|Use|Value|Def)Box(?:e(?=s))? \1 UnitBox/UnitBoxes $\rightarrow$ Unit/Units
All Soot Unit Stmt Unit/Units $\rightarrow$ Stmt/Stmts
All Soot BodyTransformer BodyInterceptor -
All SootUp BasicBlock Block -
[…]
All SpotBugs spotbugsTestCases findbugsTestCases -
Normalizing API docstrings and comments: API docstrings and inline comments may contain content that is irrelevant to code functionality, so we normalize them before applying the matching rules. For docstrings, we remove inline tags and HTML syntax, comparing only the description text (e.g., {@link Path} becomes Path). We also remove URLs and TODOs from docstrings and comments. Moreover, we convert abbreviations in natural language to their full forms to eliminate irrelevant differences (e.g., “doesn’t” is converted to “does not”).
Tokenization: We tokenize method details for similarity measurement by replacing punctuation with whitespace and splitting the text into words. We then split camel case words into individual tokens and convert all tokens to lowercase.
4.3 Method-level Post-Filtering
To ensure high-confidence mappings, we apply method-level post-filtering using a Semantic Alignment Score (SAS), which quantifies semantic similarity between method pairs and allows us to filter out candidates below a predefined threshold $thres_{SAS}$.
Semantic Alignment Score Computation: SAS quantifies semantic similarity between a method pair ($m_{1}$, $m_{2}$) from two projects. It compares class names, method names, API docstrings, and inline comments, and comprises three components: ① Class similarity (simClasssimClass) based on class names and docstrings; ② Method header similarity (simMethodHeadersimMethodHeader) based on method names, return type, parameter names and types; ③ Optional detail similarity (simOptionalsimOptional) based on names and types of local variable, method docstrings, and inline comments.
Each component is computed using the longest common subsequence (LCS) between token sequences, which has been shown effective for comparing code and documentation (Koznov et al., 2024; Roy and Cordy, 2008): $\text{sim}(S_{1},S_{2})=\frac{2\cdot\text{lcs}(S_{1},S_{2})}{n+m}$, where $S_{1}$ and $S_{2}$ have $n$ and $m$ tokens, respectively, and $lcs(S_{1},S_{2})$ is the length of their longest common subsequence. We compute LCS-based similarity for all relevant method details (simClassNamesimClassName, simClassDocsimClassDoc, simMethodNamesimMethodName, simReturnTypesimReturnType, simParamsimParam, simLocalVarsimLocalVar, simMethodDocsimMethodDoc, simCommentsimComment) and aggregate them into SAS components:
$\left\{\begin{aligned} simClass&=simClassName+(1-simClassName)\cdot simClassDoc,\\ simMethodHeader&=\delta\cdot simMethodName+\eta\cdot simReturnType+\phi\cdot simParam,\\ simOptional&=\text{average}(simLocalVar,simMethodDoc,simComment)\end{aligned}\right$.
The overall SAS is a weighted sum of these components:
$SAS=\alpha\cdot simClass+\beta\cdot simMethodHeader+\theta\cdot simOptional$
With hyperparameters $\alpha,\beta,\theta,\delta,\eta,\phi$ controlling the contribution of each part. These were tuned via grid search on labeled samples from Soot/SootUp (Table 5) to maximize true positives (TPs) among the top-$K$ most similar method pairs ($K$ being the number of TPs in the training set). When multiple combinations yielded the same TP count, we select the one maximizing the minimum hyperparameter to ensure balanced contributions. The resulting optimal values are 0.5, 0.25, 0.25, 0.5, 0.35, and 0.15, respectively.
Implementation. During data extraction, we use JavaParser (JavaParser, 2024) to analyze Java files. In preprocessing, we apply custom rules (Table 3) to standardize equivalent identifiers for Soot/SootUp and FindBugs/SpotBugs, prioritizing renaming complex identifiers to simpler ones to reduce incorrect replacements in camel-case identifiers. For SAS computation, we use the Python package pandarallel to parallelize execution.
5 Evaluation
Our evaluation aims to address the following research questions:
  • RQ1: How effective are prior code clone detection tools in identifying code and test reuse? This RQ examines how well prior clone detectors support method-level code mappings across redesigned projects, and the challenges in applying them in such scenarios.
  • RQ2: To what extent does our approach succeed in filtering irrelevant method clones? This RQ evaluates the overall effectiveness of our proposed approach in filtering irrelevant clones and improving the accuracy of the resulting code mappings.
  • RQ3: How well does each individual rule perform in filtering method clones? This RQ analyzes the contribution of each rule in our approach through an ablation study.
Clone Detector Configuration and Setup. We select four baseline clone detectors: NiCad, DeepSim, CCStokener, and an LLM-based detector (Section 2.2). Following prior studies (Van Bladel and Demeyer, 2020; van Bladel and Demeyer, 2021), we adopt a minimum clone size of 5 LOC for all tools. NiCad and CCStokener use default similarity thresholds of 0.7 and 0.6, respectively. For DeepSim, we use the version trained on Google Code Jam data rather than BigCloneBench, which has been shown to be problematic for learning code similarity (Krinke and Ragkhitwetsagul, 2022). For the LLM-based detector, we use the latest open-weight GPT-OSS-120B model because it is comparable to OpenAI’s o4-mini model (OpenAI, 2025), running on an NVIDIA H100 GPU. For all other tool configurations, we reuse their default values.
5.0.1 Dataset Construction
We evaluated baseline tools and our approach on two redesign pairs Soot/SootUp and FindBugs/SpotBugs. Table 4 presents detected clones, with “Orig” showing the initial count.
Table 4: Method pairs detected by each tool before and after post-filtering (SAS $\geq 0.5$)
Code Type Pair NiCad CCStokener DeepSim GPT-OSS-120B Orig Filt Out (%) Orig Filt Out (%) Orig Filt Out (%) Orig Filt Out (%) Production Soot/SootUp 512 182 64.45 1,518 295 80.57 1,961,526 1,159 99.94 190 139 26.84 FindBugs/SpotBugs 2,049 2,024 1.22 4,049 3,112 23.14 1,200,747 11,658 99.03 4,144 3,987 3.79 Test Soot/SootUp 2 0 100 600 4 99.33 130,614 715 99.45 9 0 100 FindBugs/SpotBugs 70 70 0 185 182 1.62 2,840 58 97.96 115 112 2.61 Average Reduction Rate 41.42 51.17 99.10 33.31
Table 5: Labeling results for method pairs in sample set $D$
CodeType Pair Total NonClones Genuine Clones Code Mappings Total T1 T2 T3 T4 Total T1 (%) 1 T2 (%) T3 (%) T4 (%) Production Soot/SootUp 539 356 183 47 50 56 30 145 100.00 58.00 76.79 86.67 FindBugs/SpotBugs 535 194 341 206 34 95 6 299 100.00 44.12 78.95 50.00 Test Soot/SootUp 209 209 0 0 0 0 0 0 0.00 0.00 0.00 0.00 FindBugs/SpotBugs 466 245 221 57 28 132 4 118 100.00 82.14 25.76 100.00 1 Columns such as “T1 (%)” indicate the percentage of Type-1 genuine clones that are code mappings.
Ground Truth Dataset. As manually examining all clone pairs is impractical, we built a ground truth dataset $D$ through manual inspection, comprising four subsets from production and test code for each redesign pair. We sampled detected clones evenly from each baseline to maintain representativeness, forming $D=NiCad\cup DeepSim\cup CCStokener\cup\text{GPT-OSS-120B}$. We initially targeted around 400 samples per set to achieve a 95% confidence level with a 5% margin of error, though the actual number varies across sets due to differences in available clone candidates. To mitigate overfitting, we excluded methods added by our ported patches and tests, as our rules were derived from the case study and some PRs have already been merged into the latest Soot/SootUp versions. This process resulted in 1,749 labeled samples in total. Two authors independently reviewed all sampled method pairs and met to resolve disagreements to ensure consistent clone labels.
Table 5 presents the number of GCs and CMs in the dataset $D$, along with the clone type distribution. Columns “T1” to “T4” list the counts of Type-1 to Type-4 clones, while “T1 (%)” to “T4 (%)” give the percentages of CMs among GCs of each type. No GCs or CMs appear in Soot/SootUp test samples. Both GCs and CMs are more prevalent in production than in test code. In production, Soot/SootUp shows a balanced distribution across Type-1 to Type-4, whereas Type-1 clones dominate in FindBugs/SpotBugs, reflecting heavier redesign in SootUp versus minor edits in SpotBugs. Most GCs are CMs (145/183 in Soot/SootUp and 299/341 in FindBugs/SpotBugs), supporting the applicability of clone detectors for identifying CMs in redesigned projects. In contrast, Type-3 clones dominate FindBugs/SpotBugs test code, with only 25% being CMs. The remaining Type-3 are not CMs where most appear to be similar due to common structures used in test harness. For example, tests in Figure 8 are flagged as clones by all baselines, but their test method names and variable names indicate different functionalities. Such irrelevant clones are confusing from a code migration perspective, as mappings for both methods already exist, highlighting the need to remove them. Our manual analysis found numerous non-clones, with three notable characteristics: (1) methods with low similarity in identifier names and functionality, (2) short methods with high syntactic similarity (e.g., null check in Figure 9), and (3) methods labeled as clones by learning-based tools but substantially differs in size and content. This shows the importance of filtering non-clones. Considering identifier name similarity can help filter both irrelevant clones and non-clones.
1- public void testArrayOfPrimitiveIsSubtypeOfObject() throws Throwable { 2+ void testStringSubtypeOfObject() throws Throwable { 3 executeFindBugsTest(new RunnableWithExceptions() { 4 Override 5- public void run() throws Exception { 6+ public void run() throws Throwable { 7 Subtypes2 test = getSubtypes2(); 8- assertTrue(test.isSubtype(typeArrayInt,typeObject));}});} 9+ Assertions.assertTrue(test.isSubtype(typeString,typeObject));}});}
Figure 8: An example of irrelevant clones that are not code mappings in test code of FindBugs/SpotBugs
1public OnFlyCallGraph getCICallGraph() {
2 if (cicg == null) {
3 constructCallGraph(); }
4 return cicg; }
(a) SootUp method in CallGraphBuilder.java
1public List<SootClass> getExceptions(){
2 if (exceptions == null) {
3 exceptions=new ArrayList<SootClass>();}
4 return exceptions; }
(b) Soot method in SootMethod.java
Figure 9: A non-clone pair identified by NiCad due to high syntactic similarity
Finding 4: Test harness within test code leads to more irrelevant clones than in production code. The challenge of filtering non-clones and irrelevant clones in code mappings can be addressed by considering identifier name similarity.
Bugs Detected in Code Mappings. While inspecting code mappings, we found three clones (SootUp, 2024b, c, d) with inconsistent behavior, suggesting potential bugs during redesign. Although CCStokener and DeepSim can detect them, they are buried among 1,518 and 1,961,526 detected clones, like finding a needle in a haystack. Our enhanced rules help prioritize these code mappings for inspection. For example, a Soot method in Figure 10(a) performs extra checks for an Int-like type, while its rewritten Type-4 clone in SootUp (Figure 10(b)) returns UnknownType. The highlighted condition should be negated. We also found clones missing a null check or an instanceof check. These issues were reported with fixes to SootUp and accepted. These examples show the importance of carefully examining changes in reused code to avoid bugs.
Finding 5: Our manual analysis of code porting found three instances of inconsistent behavior, suggesting that analyzing code mappings can help reveal porting-related bugs.
1public Type getType() {
2 if(isIntLikeType(op2Box.getValue().getType())) {
3 final Type t1 = op1Box.getValue().getType();
4 if (isIntLikeType(t1)) {
5 return IntType.v(); }
6 final LongType tyLong = LongType.v();
7 if (tyLong.equals(t1)) {
8 return LongType.v(); }}
9 return UnknownType.v(); }
(a) Soot method in JUshrExpr.java
1public Type getType() {
2 Value op1 = getOp1();
3 Value op2 = getOp2();
4 if (Type.isIntLikeType(op2.getType())) {
5 return UnknownType.getInstance(); }
6 if (Type.isIntLikeType(op1.getType())) {
7 return PrimitiveType.getInt(); }
8 if (op1.getType().equals(PrimitiveType.getLong())) {
9 return PrimitiveType.getLong(); }
10 return UnknownType.getInstance(); }
(b) The clone in JUshrExpr.java in SootUp
Figure 10: An example of a bug in correctly mapped clones
5.0.2 Evaluation Metrics
We evaluate baseline and our approach on dataset $D$ using the following metrics. In our context, true positives (TP) are genuine clones or code mappings in $D$ correctly detected, while false positives (FP) are non-clones or non-mappings incorrectly detected. True negatives (TN) are non-clones or non-mappings within $D$ correctly ignored, while false negatives (FNFN) are genuine clones or code mappings in $D$ missed by a detector. Method pairs outside $D$ are excluded from all calculations.
FPR: False Positive Rate quantifies the fraction of negative instances incorrectly classified as positive. It is calculated as $FPR=\frac{FP}{FP+TN}$.
Precision: Precision measures the fraction of correctly predicted positive instances among all instances predicted as positive, and its formula is $Precision=\frac{TP}{TP+FP}$.
Recall: Recall computes the fraction of actual positive instances that are correctly detected. It is calculated as $\frac{TP}{TP+FN}$.
Average F1-Score: The F1-score denotes the harmonic mean of precision and recall ($F1=\frac{2\cdot Precision\cdot Recall}{Precision+Recall}$). We measures the performance on the positive class (i.e., genuine clones or code mappings) but ignores the negative class (non-clones or non-mappings). To better measure the overall effectiveness, we follow prior work (Suh et al., 2025) and compute two F1 variants (denoted as $F1_{GCs/CMs}$ and $F1_{non-clones/non-mappings}$) by alternately treating the positive class as GCs/CMs or non-clones/non-mappings. The average F1-score is then calculated as the mean of these two values: $Average\ F1=\frac{F1_{GCs/CMs}+F1_{non-clones/non-mappings}}{2}$.
5.1 RQ1: Efficacy of Prior Clone Detectors
Table 6: Impact of our post-filtering on genuine clones and code mappings on sample set $D$
CodeType Pair Tool FPR 1 Precision Recall Average F1-Score Genuine Clones Prod Soot/SootUp 2 CCStokener 0.35/0.03/-0.32 0.44/0.89/0.45 0.53/0.51/-0.02 0.58/0.76/0.18 DeepSim 0.47/0.01/-0.46 0.31/0.94/0.63 0.42/0.40/-0.02 0.47/0.71/0.24 GPT-OSS-120B 0.10/0.01/-0.09 0.78/0.98/0.20 0.70/0.66/-0.04 0.81/0.85/0.04 NiCad 0.19/0.02/-0.17 0.63/0.95/0.32 0.63/0.63/0.00 0.72/0.83/0.11 FindBugs/SpotBugs 3 CCStokener 0.26/0.06/-0.20 0.84/0.96/0.12 0.80/0.79/-0.01 0.77/0.84/0.07 DeepSim 0.75/0.03/-0.72 0.51/0.96/0.45 0.44/0.43/-0.01 0.34/0.63/0.29 GPT-OSS-120B 0.05/0.01/-0.04 0.97/1.00/0.03 0.87/0.86/-0.01 0.90/0.90/0.00 NiCad 0.01/0.00/-0.01 0.99/1.00/0.01 0.74/0.72/-0.02 0.83/0.82/-0.01 Test Soot/SootUp 4 CCStokener 0.05/0.00/-0.05 0.00/0.00/0.00 0.00/0.00/0.00 0.49/0.50/0.01 DeepSim 0.48/0.00/-0.48 0.00/0.00/0.00 0.00/0.00/0.00 0.34/0.50/0.16 GPT-OSS-120B 0.04/0.00/-0.04 0.00/0.00/0.00 0.00/0.00/0.00 0.49/0.50/0.01 NiCad 0.01/0.00/-0.01 0.00/0.00/0.00 0.00/0.00/0.00 0.50/0.50/0.00 FindBugs/SpotBugs CCStokener 0.09/0.07/-0.02 0.88/0.91/0.03 0.74/0.74/0.00 0.83/0.84/0.01 DeepSim 0.87/0.01/-0.86 0.13/0.91/0.78 0.14/0.14/0.00 0.14/0.48/0.34 GPT-OSS-120B 0.03/0.00/-0.03 0.94/1.00/0.06 0.49/0.48/-0.01 0.72/0.73/0.01 NiCad 0.00/0.00/0.00 1.00/1.00/0.00 0.32/0.32/0.00 0.62/0.62/0.00 Code Mappings Prod Soot/SootUp CCStokener 0.35/0.03/-0.32 0.38/0.86/0.48 0.59/0.54/-0.05 0.59/0.79/0.20 DeepSim 0.45/0.02/-0.43 0.26/0.89/0.63 0.44/0.40/-0.04 0.48/0.72/0.24 GPT-OSS-120B 0.14/0.03/-0.11 0.67/0.89/0.22 0.77/0.70/-0.07 0.80/0.86/0.06 NiCad 0.23/0.06/-0.17 0.50/0.79/0.29 0.63/0.62/-0.01 0.68/0.80/0.12 FindBugs/SpotBugs CCStokener 0.33/0.01/-0.32 0.76/0.99/0.23 0.83/0.81/-0.02 0.75/0.89/0.14 DeepSim 0.66/0.00/-0.66 0.47/0.99/0.52 0.47/0.45/-0.02 0.41/0.68/0.27 GPT-OSS-120B 0.10/0.00/-0.10 0.92/1.00/0.08 0.94/0.92/-0.02 0.92/0.95/0.03 NiCad 0.09/0.01/-0.08 0.92/0.99/0.07 0.78/0.76/-0.02 0.84/0.86/0.02 Test Soot/SootUp CCStokener 0.05/0.00/-0.05 0.00/0.00/0.00 0.00/0.00/0.00 0.49/0.50/0.01 DeepSim 0.48/0.00/-0.48 0.00/0.00/0.00 0.00/0.00/0.00 0.34/0.50/0.16 GPT-OSS-120B 0.04/0.00/-0.04 0.00/0.00/0.00 0.00/0.00/0.00 0.49/0.50/0.01 NiCad 0.01/0.00/-0.01 0.00/0.00/0.00 0.00/0.00/0.00 0.50/0.50/0.00 FindBugs/SpotBugs CCStokener 0.34/0.04/-0.30 0.35/0.81/0.46 0.55/0.53/-0.02 0.58/0.77/0.19 DeepSim 0.64/0.00/-0.64 0.09/0.95/0.86 0.19/0.17/-0.02 0.28/0.58/0.30 GPT-OSS-120B 0.03/0.00/-0.03 0.90/0.99/0.09 0.88/0.84/-0.04 0.93/0.94/0.01 NiCad 0.10/0.01/-0.09 0.51/0.90/0.39 0.31/0.30/-0.01 0.61/0.67/0.06 1 For Soot/SootUp, similarity thresholds are 0.5 for genuine clones and 0.6 for code mappings. 2 For FindBugs/SpotBugs, similarity thresholds are 0.6 for genuine clones and 0.8 for code mappings. 3 The test samples for Soot/SootUp contain no genuine clones or code mappings, capping the average F1-score at 0.5.
Table 6 presents the evaluation results of selected tools on dataset $D$ for detecting GCs and CMs. Each value follows “X/Y/Z” format, where X is the original tool result, Y the filtered result, and Z the improvement (YXY-X). Note that for test code of Soot/SootUp, the average F1-scores remains non-zero despite zero precision and recall on the positive class (GCs/CMs) because the average F1-score incorporates the F1-scores across both positive and negative classes (non-clones/non-mappings).
We observe that prior tools face limitations in detecting GCs and CMs, especially for extensively redesigned projects. For GC detection, both CCSTokener and DeepSim tend to produce many FPs, with FPRs of 0.26 and 0.75 on Soot/SootUp production code. DeepSim performs worse than other baselines across precision, recall, and average F1-score, while CCSTokener performs better on FindBugs/SpotBugs than on Soot/SootUp (precision 0.84 and 0.44, respectively). This difference likely stems from the more extensive redesign of Soot/SootUp, which introduces more modifications and makes clone detection harder. GPT-OSS-120B and NiCad generally achieve the best overall performance across metrics. However, both suffer performance drops in highly redesigned projects, e.g., their FPRs increase from 0.05 and 0.01 in FindBugs/SpotBugs to 0.10 and 0.19 in Soot/SootUp, while precision and recall decrease. These observations indicate the need for more robust strategies to reduce FPs under extensive redesigns.
Compared to GC detection, all tools reported lower precision for CM detection as some detected pairs are genuine clones but not code mappings. For example, NiCad’s precision drops from 1.00 to 0.51 on FindBugs/SpotBugs test code, as many tests share similar API usage and structure but differ in input and expected output, targeting different functionalities. In contrast, recall increases because CMs, as a subset of GCs, have a smaller denominator. This underscore the importance of distinguishing CMs from GCs to build accurate code mappings in redesigned projects.
Finding 6: Compared to test code, prior clone detection tools are more effective at identifying reusable production code. These tools need to be adapted for more effective code mapping detection (e.g., by enhancing false positive filtering for non-clones and irrelevant clones).
5.2 RQ2: Filtering Effectiveness
5.2.1 Pre-filtering Effectiveness
We applied class-level pre-filtering (Section 4.1) to enable repository-level processing specifically for GPT-OSS-120B. This step reduces the total pairs from 12.3M to 112K (99% filtered) for Soot/SootUp and from 15.6M to 273K (98% filtered) for FindBugs/SpotBugs, allowing the remaining pairs to be processed in about 12 and 29 hours, respectively.
5.2.2 Post-filtering Effectiveness
To evaluate our post-filtering approach, we computed SAS for method pairs in dataset $D$, and determined the SAS threshold $thres_{SAS}$ via Average F1-score. Figure 11 shows GC and CM detection results in Soot/SootUp across $thres_{SAS}$, using SAS alone and with baseline tools. The optimal SAS threshold $thres_{SAS}$ is the same for production and test code, with code mappings generally requiring higher values than genuine clones. Accordingly, we set $thres_{SAS}$ for Soot/SootUp to 0.5 (GC) and 0.6 (CM). We observed a similar pattern for FindBugs/SpotBugs, and thus set $thres_{SAS}$ to 0.6 (GC) and 0.8 (CM).
(a) GC Detection on Soot/SootUp
(b) CM Detection on Soot/SootUp
Figure 11: SAS and baseline integration across thresholds
Table 6 shows the impact of our filtering approach. It effectively reduces FPs, achieving very low FPRs (maximum 0.07) and improving precision for all baselines on both GC and CM detection (e.g., DeepSim’s FPR drops by 0.64 and precision rises by 0.86 on FindBugs/SpotBugs test code). Although recall decreases slightly (mostly within –0.02), average F1-scores generally improve, indicating the efficacy of our approach. GPT-OSS-120B (top baseline) also shows solid gains, increasing its precision on Soot/SootUp production code from 0.78 to 0.98, with only a minor recall drop (–0.04). For GPT-OSS-120B, although applying our rules in the post-filtering stage may appear to yield moderate improvement for the average F1-score, this may be due to the fact that substantial pruning has occurred at the pre-filtering stage (which enables GPT-based approach to focus on a selected set of clones). Moreover, our filtering helps distinguish CMs from GCs, e.g., NiCad’s precision for CM detection in FindBugs/SpotBugs test code rises from 0.51 to 0.90 after integration, compared to its previous 1.00 precision for GC detection.
Time Overhead. We measured the execution time of our post-filtering approach (including preprocessing and SAS computation) on dataset set $D$, which was sampled from the union of clones detected by all tools. The dataset contains 748 clones for Soot/SootUp and 1001 clones for FindBugs/SpotBugs. Experiments were run on an Apple M3 machine (8-core CPU, 20GB RAM) for five runs. The average execution times of 2.75s for Soot/SootUp and 3.52s for FindBugs/SpotBugs show that our rules introduce minimal overhead.
Post-filtering on Raw Detection Results. To evaluate the practical impact of our post-filtering approach beyond labeled datasets, we applied a SASSAS threshold of 0.5, the minimum optimal value observed in previous analysis, to the raw results detected by the selected techniques. Table 4 shows the filtered counts (“Filt”) and the percentage of removed pairs (“Out (%)”). We observed that DeepSim identified millions of production code clones, likely due to the large search space (e.g., comparing all pairs among 3,896 FindBugs and 3,970 SpotBugs methods in production code). The sheer volume makes manual verification impractical, highlighting the need for effective filtering. By filtering low-similarity clones, our rules help reduce the manual inspection cost of analyzing these clones. The average reduction rate is 33–99%.
Finding 7: Our rules improve clone detection for redesigned projects: pre-filtering prunes over 98% of method pairs for GPT-based detectors, enabling repository-scale analysis, while post-filtering reduces false positives, boosting precision by up to 86% and cutting manual inspection by 33–99%.
5.3 RQ3: Efficacy of each rule
To assess the impact of individual rules in Table 2, we exclude a rule by removing the corresponding code information during SAS computation (Equation 1), resulting in five distinct settings: ALL (all rules applied), EXR1 (exclude R1 by disabling custom renaming rules in preprocessing), EXR2 (exclude R2 with $simLocalVar=simMethodHeader=0$), EXR3 (exclude R3 with $simMethodDoc=simClassDoc=0$), and EXR4 (exclude R4 with $simComment=0$).
For each setting, we compute SAS for method pairs in dataset $D$, rank them, and calculate average F1-scores across SAS thresholds. Figure 12 shows the effect on CM detection (GC detection exhibits similar trends), with identifier similarity having the largest impact: excluding it in $EXR2$ noticeably reduces the average F1. To better understand each rule’s contribution, we track method pairs where SAS changes after exclusion, and record the number of affected pairs, maximum SAS change and maximum rank change (Table 7). R2 has the largest effect on ranking, with max SAS and rank changes of 0.83 and 141, as identifiers are more common than API documentation or comments. Other rules affect fewer methods but capture complementary semantic cues. For example, R1 affects few methods but still produces considerable SAS and rank changes (up to 0.23 and –63), indicating that even local renaming impacts semantic alignment. Together, these rules yield a more accurate semantic similarity and help filter otherwise undetected FPs.
Finding 8: Among all implemented rules, the “R2: Similar Identifier Name Matching” rule is the most effective one, while the other rules enhance similarity measurement by leveraging different method information.
(a) CM Detection on Soot/SootUp
(b) CM Detection on FindBugs/SpotBugs
Figure 12: Impact of rule exclusion across SAS thresholds
Table 7: Method pairs affected by each rule in dataset $D$
Soot &SootUp/ R1 R2 R3 R4 FindBugs &SpotBugs Prod Test Prod Test Prod Test Prod Test # Affected 63 5 441 202 142 4 55 15 Max SAS Change 0.23 0.06 0.83 0.62 0.19 0.01 0.08 0.04 Max Rank Change -69 -19 170 80 59 -4 -28 -15 # Affected 3 0 487 466 139 4 133 117 Max SAS Change -0.00 0.00 0.83 0.83 0.08 0.08 0.08 0.08 Max Rank Change 2 0 122 141 27 -5 -16 24
6 Threats to Validity
External. Our approach improves efficacy across four evaluated tools. As the efficacy may vary with other tools, we selected diverse clone detection techniques to enhance representation. Our approach is effective on the evaluated samples but may vary on other datasets. To mitigate this, we evaluated it on a fairly large dataset of over 385 samples to ensure a 95% confidence level. Moreover, our results may not generalize beyond Soot/SootUp and FindBugs/SpotBugs. To enhance general applicability, we (1) derived our rules from the bidirectional study on Soot/SootUp and evaluated on two project pairs, and (2) reported any faulty clone pairs to the developers.
Internal. We manually compared each clone pair for correctness. To reduce potential bias, two authors independently verified the labels and resolved any conflicts through discussion.
7 Related Work
Studies. Prior work has examined multiple versions of a software project (Shariffdeen et al., 2021; Liew et al., 2017), e.g., developers’ experiences in redesigning their own projects (Dorman and Arnaud, 2001; Karakaya et al., 2024). Our study can be seen as a special case of N-version programming, e.g., a study examined N-version programming where two research teams add floating-point support for KLEE independently (Liew et al., 2017). Another study focuses on patch backporting in the Linux kernel (Shariffdeen et al., 2021). Our study differs from prior studies in two key aspects: (1) we study redesigned projects instead of program versions, and (2) we focus on code and test reuse practices instead of other practices (e.g., adding new support (Liew et al., 2017) or backporting patches (Shariffdeen et al., 2021)).
Code clone detection. Although some code clone detection tools target specific domains (e.g., smart contracts (He et al., 2020), apps (Wang et al., 2015), pre-training (Ding et al., 2023), or clone synchronization (Cheng et al., 2016)), we focuse on general-purpose code clone detection tools introduced in Section $2.$ In future, it is worthwhile to study enhancing domain-specific clone detection for redesigned projects. Prior approaches typically rely on abstract representations (e.g., program dependence graphs) or code normalization to capture similarity, but often ignore identifier names and code comments, which are crucial in redesigned projects. Documentation contains valuable domain knowledge, yet many tools ignore it due to the ambiguity of natural language. Several studies have confirmed its usefulness for code clone detection (Kuttal and Ghosh, 2020; Gupta and Goyal, 2021), e.g., Kuttal et al (Kuttal and Ghosh, 2020) use Latent Dirichlet Allocation (LDA) at file level combined with PDGs for function-level detection, while Gupta and Goyal (Gupta and Goyal, 2021) extract method documentation and apply Latent Semantic Indexing (LSI) with various similarity measures to identify concept-level clones. In contrast, our approach leverages near-duplicate documentation in redesigned projects using a simpler LCS-based method (Koznov et al., 2024), and combines it with identifier information. Moreover, our rules can be applied to prior code clone detection tools for further effectiveness enhancement.
Identifier Similarity. Prior work shows that combining identifier similarity with structural information is effective for clone detection. Marcus et al. (Marcus and Maletic, 2001) and CLAN (McMillan et al., 2012) uses LSI to capture conceptual similarity, using code comments and identifiers or API calls from common libraries. Higo et al. (Higo and Kusumoto, 2014) shows that combining vocabulary similarity with structural similarity reduces false positives, but vocabulary-based measures are less effective for method pairs within the same file, where methods often share fields. Our SAS is related to vocabulary similarity but is tailored to redesigned projects. Specifically, we (1) model identifiers in a finer-grained manner, e.g., decomposing signature similarity into return type, method name, and parameters; (2) preserve all identifier tokens instead of filtering by part of speech or stop words, since identifiers are short and often reused or slightly modified during redesign; and (3) use LCS rather than Jaccard similarity to retain token order, which is usually preserved under local modifications and helps mitigate limitations of vocabulary similarity for intra-file methods.
Code reuse. Researchers have explored code reuse to avoid re-implementing functionality (Gharehyazie et al., 2017; Wang et al., 2016; Haefliger et al., 2008), including techniques for reusing low-level code (De Sutter et al., 2002; Mehta et al., 2023). Unlike prior work, we (1) propose rules to improve clone detection and establish code mappings between redesign and original projects, and (2) focus on redesigned projects rather than general cross-project clones.
8 Lessons Learned
We discuss the lesson learned and implications based on our study.
[…]
Beyond static analyzers. Although our study focuses on static analyzers (a challenging cross-repo scenario where the redesign needs often arise), we believe that all findings are generally applicable beyond static analyzers as our action research translates these findings to real actions by making open-source contributions. Moreover, as most of our proposed rules are general (except for R1 where users need to add customized replacement rules), they can be directly applied to enhance existing clone detection techniques to uncover the code mappings between the redesigned projects.
Automated change tracking in redesigned projects. As the migration to SootUp is still an ongoing effort, developers must maintain both versions until SootUp becomes “feature-complete”. When both projects evolve, it leads to divergence in code and tests, scattering related changes across multiple artifacts (Finding 1). To find all relevant changes, we need to conduct both keyword and label searches across source code, commits, issues/PRs, and documentation,indicating the need for automated tools to track relevant changes during redesign. One possible solution is adding a GitHub feature (GitHubCommunity, 2022) to support linking a PR or issue to multiple projects, but it only partially solves the problem as it misses relevant discussions in documentation. Hence, tools that can track relevant changes, focusing on the mapping between the changes between the original and redesigned projects, are a worthwhile direction to explore. Our rules, which can be applied on top of prior code clone detection tools can help reestablish these links.
Technical debts as deferred code reuse. Our study revealed that technical debts exist in TODO comments (C1) as a hidden form of deferred code reuse. Adapting the code within the TODOs often required substantial efforts as they are partially completed and outdated code. In future, it is worthwhile to design code generation techniques that can automatically reuse and adapt code.
Test reuse for static analyzers. Our study reveals missed opportunities for backporting tests between the two static analysis projects, as developers often rewrite tests instead of reusing them, increasing maintenance burden. While several automated testing approaches for static analyzers (Mordahl et al., 2023; Zhang et al., 2023a; He et al., 2024; Zhang et al., 2024a, b; Kaindlstorfer et al., 2024; Li et al., 2023) have been proposed, they do not naturally support test reuse. Future work can enhance these techniques to better facilitate test reuse.
Importance of naming information. Our evaluation shows that SAS that captures semantic information embedded within names and documentations helps improve clone detection for redesigned projects. Among all information, Finding 8 shows that identifier names is the effective in finding code mappings. This highlights the importance of naming information, and provides empirical evidence for future research on identifier renaming in redesigned projects.
9 Conclusion
This paper investigates software reuse in cross-repository redesigns using an action research–driven, bidirectional study of the ongoing Soot/SootUp redesign case. Our study shows that redesign is non-linear and often requires bidirectional reuse, while tests and residual bugs are frequently overlooked during migration. We identify tracking corresponding code and tests as the key challenge and address it by retrofitting clone detection using filtering based on Semantic Alignment Scores. Our evaluation on redesign cases (Soot/SootUp and FindBugs/SpotBugs) shows that our approach improves precision and reduces manual effort, and our open-source contributions demonstrate practical impact. These results provide actionable insights and tools to better support software reuse in ongoing redesign projects.
Data Availability Our implementation, data, and scripts are available at https://doi.org/10.5281/zenodo.19324267, with the analyzed project versions being Soot 4.5.0, SootUp 1.3.0, SpotBugs 4.5.0, and FindBugs 3.0.1.